OpenStudy (anonymous):

when p(x)=x^2 + ax -b is divided by x-1, the remainder is -6. when p(x) is divided by x+2. the remainder is 3. find the values of a and b.

OpenStudy (anonymous):

@phi pls help

terenzreignz (terenzreignz):

Simple... if f(x) is any polynomial, the remainder when f(x) is divided by x-c is equal to f(c)

OpenStudy (anonymous):

how?

terenzreignz (terenzreignz):

Let me do the first bit for you... In the case of x - 1, c = 1, right?

OpenStudy (anonymous):

yup

terenzreignz (terenzreignz):

Right... Now, to get the remainder when p(x) is divided by x-1 it is simply p(1) Now what is p(1) equal to?

OpenStudy (anonymous):

x=1

terenzreignz (terenzreignz):

I mean, plug in x = 1 in p(x) Replace all the x's with 1.

OpenStudy (anonymous):

I don't get it, sorry

terenzreignz (terenzreignz):

@derek11111 I'm going out on a limb here and assuming that you're given the same question.... Get over here :P

OpenStudy (anonymous):

http://www.purplemath.com/modules/remaindr.htm

OpenStudy (anonymous):

xd, ahahah

OpenStudy (anonymous):

HAHAHHAA

OpenStudy (anonymous):

\[p(x) / (x-a) = q(x) + r(x) \] where r(x) is the remainder

OpenStudy (anonymous):

so \[p(x) = q(x)(x-a) + r(x)\]

OpenStudy (anonymous):

now if you plug in p(x) x=a you get the remainder.

terenzreignz (terenzreignz):

Now listen up, the two of you -.- To get the remainder of a polynomial p(x) when it's divided by x - c You simply evaluate that polynomial at x = c p(c) In this case, we are dividing by x - 1 So c = 1 Now... \[\Large p(\color{red}x)=\color{red}x^2 + a\color{red}x - b\]

terenzreignz (terenzreignz):

So... \[\Large p(\color{red}1) = \color{red}1^2 + a(\color{red}1) -b\]

OpenStudy (anonymous):

so for your case, p(x)=x^2 + ax -b divided by x-1 gives remainder of -6 so p(1) = -6

OpenStudy (anonymous):

but look at the website or my first responses so youll understand why this is true.

OpenStudy (anonymous):

what is the value of a and b? o.O?

OpenStudy (anonymous):

you will get two equations : p(1) = -6 and p(-2) = 3

OpenStudy (anonymous):

using them youll find a and b.

OpenStudy (anonymous):

more important is to understand why it is working though.

OpenStudy (anonymous):

can you please tell me how to find a and b using those 2

OpenStudy (anonymous):

p(1) = 1+a-b so since p(1) = -6 1+a-b=-6 a-b=-7 p(-2) = 4-2a-b so since p(-2) = 3 4-2a-b=3 -2a-b=-1 you have two equations : a-b=-7 -2a-b=-1 you can subtract one from the other and find b. then plug b into one of them and find a.

OpenStudy (anonymous):

how to subtract a-b= -7 from -2a-b= -1?

OpenStudy (anonymous):

subtract each side : -2a - b -(a-b) = -1 - (-7) -2a - a = 6 -3a = 6 a = -2

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

when you find b, verify your answer

OpenStudy (anonymous):

i might done some calculation mistake

OpenStudy (anonymous):

where?

OpenStudy (anonymous):

i dont say i did.. i might.. that is why you better verify your answer (check that p(1) really comes out as -6 )

OpenStudy (anonymous):

b= -3 am I right?

OpenStudy (anonymous):

no, if a =-2 then b = 5

OpenStudy (anonymous):

now I know how to do it thanks :)